OpenStudy (flyin210):

What is the force of gravity between Earth (mass = 6.0 × 1024 kilograms) and Jupiter (mass = 1.901 × 1027 kilograms)? The distance between the two planets is about 7.5 × 1011 meters. 2.027 × 1017 newtons 7.10 × 1017 newtons 1.352 × 1018 newtons 1.014 × 1019 newtons 1.54 × 1019 newtons

OpenStudy (flyin210):

pls halp

OpenStudy (irishboy123):

\(F = \dfrac{\color{red}{G} \color{blue}{m_J }\color{green}{m_E}}{r^2}\) \( = \dfrac{\color{red}{(6.67408 × 10^{-11})}\color{blue}{(1.901 × 10^{27}) }\color{green}{(6.0 × 10^{24})}}{(7.5 × 10^{11})^2}\) so, it's C ..... !!!